OpenStudy (anonymous):

what is the derivative of 5e^(-2cos^2(x))

OpenStudy (dls):

Let \(\LARGE y=5e^{(-2\cos^2(x))}\) Basic chain rule. \[\LARGE \frac{dy}{dx}=5e^{(-2\cos^2(x))} \times (4\cos x \sin x)\]